chuan invested some money in a savings account earning 6% interest and an equal amount of money in bonds earning 8% interest. How much did he invest altogether if he earned $210 in a year?
Answer by mananth(16949)

money invested = x in each
x+x=2x
6x+8x=210*100 ( percentage converted by multiplying by 100)
14x=21000
x=21000/14
x=1500
1500@6%=90
1500@8%=120
Total = 210